After changing the material of the Balance Mass from 6061 Al to 304 Steel, and changing the thickness to 0.21" from 0.25". The CoM is now 1.11mm below the clamping point.
Koji expected a mass change of ~ 4g to move the mass to 1.1mm. The 6061 mass weighed ~1.31g and the 304 mass weighs 4.1g.
A potential issue with this is the screw used the adjust the position of these balance masses, threads through both the aluminum ring and this now 304 steel mass. A non silver plated screw could cold weld at the mass, but a silver plated screw will gall in the aluminum threads.
The current vertical distance between the CoM and the wire clamping point on the 3" Ring assembly is 0.33mm. That is the CoM is .33 mm below the clamping point of the wire. I took the clamping point to be the top edge of the wire clamp piece. see the below attachments.
I am now modifying the dumbell mechanism at the bottom of the ring to move the CoM to the target distance of 1.1mm.
I attempted a single arm actuation calibration using IR beatnote (in the directions of soCal idea for DARM calibration)
An update on recent progress in the lab towards building and testing the new FEs.
The previously reported problem with the IOPs losing sync after a few minutes (16130) was resolved through a change in BIOS settings. However, there are many required settings and it is not trivial to get these right, so I document the procedure here for future reference.
The CDS group has a document (T1300430) listing the correct settings for each type of motherboard used in aLIGO. All of the machines received from LLO contain the oldest motherboards: the Supermicro X8DTU. Quoting from the document, the BIOS must be configured to enforce the following:
• Remove hyper-threading so the CPU doesn’t try to run stuff on the idle core, as hyperthreading simulate two cores for every physical core.
• Minimize any system interrupts from hardware, such as USB and Serial Ports, that might get through to the ‘idled’ core. This is needed on the older machines.
• Prevent the computer from reducing the clock speed on any cores to ‘save power’, etc. We need to have a constant clock speed on every ‘idled’ CPU core.
I generally followed the T1300430 instructions but found a few adjustments were necessary for diskless and deterministic operation, as noted below. The procedure for configuring the FE BIOS is as follows:
After completing the BIOS setup, I rebooted the new FEs about six times each to make sure the configuration was stable (i.e., would never hang during boot).
With the timing issue resolved, I proceeded to build basic user models for c1bhd and c1sus2 for testing purposes. Each one has a simple structure where M ADC inputs are routed through IIR filters to an M x N output matrix, which forms linear signal combinations that are routed to N DAC outputs. This is shown in Attachment 1 for the c1bhd case, where the signals from a single ADC are conditioned and routed to a single 18-bit DAC. The c1sus2 case is similar; however the Contec BO modules still needed to be added to this model.
The FEs are now running two models each: the IOP model and one user model. The assigned parameters of each model are documented below.
The user models were compiled and installed following the previously documented procedure (15979). As shown in Attachment 2, all the RTS processes are now working, with the exception of the DAQ server (for which we're still awaiting hardware). Note that these models currently exist only on the cloned copy of the /opt/rtcds disk running on the test stand. The plan is to copy these models to the main 40m disk later, once the new FEs are ready to be installed.
I installed several new AA and AI chassis in the test stand to interface with the ADC and DAC cards. This includes three 16-bit AA chassis, one 16-bit AI chassis, and one 18-bit AI chassis, as pictured in Attachment 3. All of the AA/AI chassis are powered by one of the new 15V DC power strips connected to a bench supply, which is housed underneath the computers as pictured in Attachment 4.
These chassis have not yet been tested, beyond verifying that the LEDs all illuminate to indicate that power is present.
I have received the opto-isolator needed to complete the new c1auxey system. I left it sitting on the electronics bench next to the Acromag chassis.
Here is the manufacturer's wiring manual. It should be wired to the +15V chassis power and to the common return from the coil driver, following the instructions herein for NPN-style signals. Note that there are two sets of DIP switches (one on the input side and one on the output side) for selecting the mode of operation. These should all be set to "NPN" mode.
Here's an updated X ARM ALS noise budget.
We went near the MC2 area and opened the lid to inspect the GigE and analog video monitors for MC2. Looked like whatever image is coming through the viewport is split into the GigE (for beam tracking) and the analog monitor. We hooked the monitor found on the floor nearby and tweaked the analog video camera around to get a feel for how the "ghost" image of the transmission moves around. It looks like in order to try and remove this "extra spots" we would need to tweak the beam tracking BS. We will consult the beam tracking authorities and return to this.
I was preparing a short write-up / test procedure for the custom HV coil driver, when I thought of something I can't resolve. I'm probably missing some really basic physics here - but why do we not account for the shot noise from DC current flowing through the series resistor? For a 4kohm resistor, the Johnson current noise is ~2pA/rtHz. This is the target we were trying to beat with our custom designed HV bias circuit. But if there is a 1 mA DC current flowing through this resistor, the shot noise of this current is 18pA/rtHz, which is ~9 times larger than the Johnson noise of the same resistor. One could question the applicability of this formula to calculate the shot noise of a DC current through a wire-wound resistor - e.g. maybe the electron transport is not really "ballistic", and so the assumption that the electrons transported through it are independent and non-interacting isn't valid. There are some modified formulae for the shot noise through a metal resistor, which evaluates to 10pA/rtHz for the same 4kohm resistor, which is still ~5x the Johnson noise.
In the case of the HV coil driver circuit, the passive filtering stage I added at the output to filter out the excess PA95 noise unwittingly helps us - the pole at ~0.7 Hz filters the shot noise (but not the Johnson noise) such that at ~10 Hz, the Johnson noise does indeed dominate the total contribution. So, for this circuit, I think we don't have to worry about some un-budgeted noise. However, I am concerned about the fast actuation path - we were all along assuming that this path would be dominated by the Johnson noise of the 4kohm series resistor. But if we need even 1mA of current to null some DC DARM drift, then we'd have the shot noise contribution become comparable, or even dominant?
I looked through the iLIGO literature, where single-stage suspensions were being used, e.g. Rana's manifesto, but I cannot find any mention of shot noise due to DC current, so probably there is a simple explanation why - but it eludes me, at least for the moment. The iLIGO coil drivers did not have a passive filter at the output of the coil driver circuit (at least, not till this work), and there isn't any feedback gain for the DARM loop at >100 Hz (where we hope to measure squeezing) to significantly squash this noise.
Attachment #1 shows schematic topologies of the iLIGO and proposed 40m configs. It may be that I have completely misunderstood the iLIGO config and what I've drawn there is wrong. Since we are mainly interested in the noise from the resistor, I've assumed everything upstream of the final op-amp is noiseless (equivalently, we assume we can sufficiently pre-filter these noises).
Attachment #2 shows the relative magnitudes of shot noise due to a DC current, and thermal noise of the series resistor, as a function of frequency, for a few representative currents, for the slow bias path assuming a 0.7Hz corner from the 4kohm/3uF RC filter at the output of the PA95.
Some lit review suggests that it's actually pretty hard to measure shot noise in a resistor - so I'm guessing that's what it is, the mean free path of electrons is short compared to the length of the resistor such that the assumption that electrons arrive independently and randomly isn't valid. So Ohm's law dictates and that's what sets the current noise. See, for example, pg 432 of Horowitz and Hill.
Here is our first attempt at a single-arm noise budget for ALS.
Attachment 1 shows the loop diagram we used to calculate the contribution of different noises.
Attachment 2 shows the measured noise at C1:ALS-BEATX_PHASE_FINE_OUT_HZ when XARM was locked to the main laser and Xend Green laser was locked to XARM.
All remaining chasses have been reworked and placed on the floor along the west wall in Room 104.
The test was succesful and brought back the IMC to lock point at the end.
We calculated new input matrix using same code in scripts/SUS/InMatCalc/sus_diagonalization.py . Attachment 1 shows the results.
The calculations are present in scripts/SUS/InMatCalc/MC1.
We uploaded the new MC1 input matrix at:
Unix Time = 1621963200
GPS Time = 1305998418
This was done by running python scripts/SUS/general/20210525_NewMC1Settings/uploadNewConfigIMC.py on allegra. Old IMC settings (before Paco and I started workin on 40m) can be restored by running python scripts/SUS/general/20210525_NewMC1Settings/restoreOldConfigIMC.py on allegra.
Everything looks as stable as before. We'll look into long term trends in a week to see if this helped at all.
Differential misalignment of the OMCs
40m BHD will employ two OMCs on the BHD platform. We will have two SOSs for each of the LO and AS beams. The challenge here is that the input beam must optimally couple to the OMCs simultaneously. This is not easy as we won't have independent actuators for each OMC. e.g. The alignment of the LO beam can be optimally adjusted to the OMC1, but this, in general, does not mean the beam is optimally aligned to the OMC2.
When a beam with the matched mode to an optical cavity has a misalignment, the power coupling C can be reduced from the unity as
where is the waist radius, is the divergence angle defined as , and are the beam lateral translation and rotation at the waist position.
The waist size of the OMC is 500um. Therefore = 500um and = 0.68 mrad. If we require C to be better than 0.995 according to the design requirement document (T1900761). This corresponds to (only) to be 35um and (only) to be 48urad. These numbers are quite tough to be realized without post-installation adjustment. Moreover, the OMCs themselves have individual differences in the beam axis. So no matter how we set the mechanical precision of the OMC installation, we will introduce a maximum of 1mm and ~5mrad uncertainty of the optical axis.
Suppose we adjust the incident beam to the OMC placed at the transmission side of the BHD BS. The reflected beam at the BS can be steered by picomotors. The distance from the BS to the OMC waist is 12.7" (322mm) according to the drawing.
So we can absorb the misalignment mode of (, ) = (0.322 , ). This is a bit unfortunate. 0.322m is about 1/2 of the rayleigh range. Therefore, this actuation is still angle-dominated but a bit of translation is still coupled.
If we enable to use the third picomotor on the BHD BS mount, we can introduce the translation of the beam in the horiz direction. This is not too huge therefore we still want to prepare the method to align the OMC in the horiz direction.
The difficult problem is the vertical alignment. This requires the vertical displacement of the OMC. And we will not have the option to lower the OMC. Therefore if the OMC2 is too high, we have to raise the OMC1 so that the resulting beam is aligned to the OMC2. i.e. we need to maintain the method to raise both OMCs. (... or swap the OMCs). From the images of the OMC beam spots, we'll probably be able to analyze the intracavity axes of the OMCs. So we can always place the OMC with a higher optical axis at the transmission side of the BHD BS.
We've set a free swing test to trigger at 3:30 am tomorrow for MC1. The script for tests is running on tmux session named 'freeSwingMC1' on rossa. The script will run for about 4.5 hrs and we'll correct the input matrix tomorrow from the results. If anyone wants to work during this time (3:30 am to 8:00 am), you can just kill the script by killing tmux session on rossa. ssh into rossa and type tmux kill-session -t freeSwingMC1.
We should redo the MC1 input matrix optimization and the coil balancing afterward as we did everything based on the noisy UL OSEM values.
Updated IOO.strip on Zita to show WFS2 pitch and yaw trends (C1:IOO-WFS2_PIY_OUT16 and C1:IOO-WFS2_YAW_OUT16) and changed the colors slightly to have all pitch trends in the yellow/brown band and all yaw trends in the pink/purple band.
No one says, "Here I am attaching a cool screenshot, becuz else where's the proof? Am I right or am I right?"
Mon May 24 18:10:07 2021 [Update]
After waiting for some traces to fill the screen, here is a cool screenshot (Attachment 1). At around 2:30 PM the MC unlocked, and the BS_Z (vertical) seismometer readout jumped. It has stayed like this for the whole afternoon... The MC eventually caught its lock and we even locked XARM without any issue, but something happened in the 10-30 Hz band. We will keep an eye on it during the evening...
Tue May 25 08:45:33 2021 [Update]
At approximately 02:30 UTC (so 07:30 PM yesterday) the 10-30 Hz seismic step dropped back... It lasted 5 hours, mostly causing BS motion along Z (vertical) as seen by the minute trend data in Attachment 2. Could the MM library have been shaking? Was the IFO snoring during its afternoon nap?
- High priority units: 2x 18AI / 1x 16AI / 3x 16AA
All six are reworked and on the electronics workbench. The rest should be ready by the end of the week.
The new HAM-A coil drivers have a single DB9 connector for all the binary inputs. This requires that the dewhitening switching signals from the fast system be spliced with the coil enable signals from c1auxey. There is a common return for all the binary inputs. To avoid directly connecting the grounds of the two systems, I have looked for a suitable opto-isolator for the c1auxey signals.
I best option I found is the Ocean Controls KTD-258, a 4-channel, DIN-rail-mounted opto-isolator supporting input/output voltages of up to 30 V DC. It is an active device and can be powered using the same 15 V supply as is currently powering both the Acromags and excitation. I ordered one unit to be trialed in c1auxey. If this is found to be good solution, we will order more for the upgrades of c1auxex and c1susaux, as required for compatibility with the new suspension electronics.
The plant transfer function of the pendulum in the s domain is:
Using Foton to make a plot of the TF needed and using m=40kg, w0=3Hz, and Q=50 (See attachment 1). It is easiest to enter the above filter using RPoly and saved it as Plant_V1
We went into 40m to identify where XARM PDH loop control elements are. We didn't touch anything, but this is to note we went in there twice at 10 AM and 11:10 AM.
The transfer function given in the previous post was slightly incorrect the units did not make sense the new function is:
I have attached a quick derivation below in attachment 1
DC Power Strip Assemblies delivered and stored behind the Y arm tube (Attachment 1)
I also moved the spare 1U Chassis to the same place.
11 new Satellite Amps were picked up from Downs. 7 more are coming from there. I have one spare unit I made. 1 sat amp has already been used at MC1.
We had 8 HAM-A coil drivers delivered from the assembling company. We also have two coil drivers delivered from Downs (Anchal tested)
HP HV power supply ( HP6209 ) were returned to Downs
For future reference, the new settings can be upoaded from a script in the same directory. Run python /users/anchal/20210505_IMC_Tuned_SUS_with_Gains/uploadNewConfigIMC.py from allegra.
There isn't any instruction here on how to upload the new settings
Calculation for the SOS POS/PIT/YAW resonant frequencies
- Nominal height gap between the CoM and the wire clamping point is 0.9mm (cf T970135)
- To have the similar res freq for the optic with the 3" metal sleeve is 1.0~1.1mm.
As the previous elog does not specify this number for the current configuration, we need to asses this value and the make the adjustment of the CoM height.
Fluke. Temp fluctuations are as usual, but the overall temperature is still lower. We ought to put some temperature sensors at the X & Y ends to see what's happening there too.
Looks like the fan lowered the temperature as expected. Need to get a few more days of data to see if its stabilized, or if that's just a fluke.
The vertical line at 00:00 UTC May 18 is about when I turned the fans up/on.
I want to work on the IFO this weekend, so I reverted the IMC suspension settings just now to what I know work (until the new settings are shown quantitatively to be superior). There isn't any instruction here on how to upload the new settings, so after my work, I will just restore from a burt-snapshot from before I changed settings.
In the process, I found something odd in the MC2 coil output filter banks. Attachment #1 shows what it it is today. This weird undetermined state of FM9 isn't great - I guess this flew under the radar because there isn't really any POS actuation on MC2. Where did the gain1 filter I installed go? Some foton filter file corruption? Eventually, we should migrate FM7,FM8-->FM9,FM10 but this isn't on my scope of things to do for today so I am just putting the gain1 filter back so as to have a clean FM9 switched on.
The old setting can be restored by running python3 /users/anchal/20210505_IMC_Tuned_SUS_with_Gains/restoreOldConfigIMC.py from allegra or donatella.
I wrote the values from the c1mcs burt snapshot from ~1400 Saturday May 15, at ~1600 Sunday May 16. I believe this undoes all my changes to the IMC suspension settings.
The NPRO has been off since ~1AM this morning it looks like. Is this intentional? Can I turn it back on (or at least try to)? The interlock signal we are recording doesn't report getting tripped but I think this has been the case in the past too.
After getting the go ahead from Koji, I turned the NPRO back on, following the usual procedure of diode current ramping. PMC and IMC locked. Let's see if this was a one-off or something chronic.
The PSL was too hot, so I turned on the south HEPA on the PSL. The north one was on and the south one was off (or so slow as to be inaudible and no vibration, unlike the north one). Lets watch the trend over the weekend and see if the temperature comes down and if the PMC / WFS variations get less. Fri May 14 17:46:26 2021
I believe I did the identical test with the one in [40m ELOG 15786]. The + input of PA95 was shorted to the ground to exclude the noise from the bias input. The voltage noise at TP6 was measured with +/-300V supply by two HP6209 and two Matsusada R4G360.
With R4G360, the floor level was identical and 60Hz line peaks were less. It looks like R4G360 is cheap, easier and precise to handle, and sufficiently low noise.
Koji and I did a few tests with an OSEM emulator on the satellite amplifier box used for MC1 which is housed on 1X4. This sat box unit is S2100029 D1002812 that was recently characterized by me 15803. We found that the differential output driver chip AD8672ARZ U2A section for the UL PD was not working properly and had a fluctuating offset at no input current from the PD. This was the cause of the ordeal of the morning. The chip was replaced with a new one from our stock. The preliminary test with the OSEM emulator showed that the channel has the correct DC value.
In further testing of the board, we found that the channel 8 LED driver was not working properly. Although this channel is never used in our current cable convention, it might be used later in the future. In the quest of debugging the issue there, we replaced AD8672ARZ at U1 on channel 8. This did not solve the issue. So we opened the front panel and as we flipped the board, we found that the solder blob shorted the legs of the transistor Q1 2N3904. This was replaced and the test with the LED out and GND shorted indicated that the channel is now properly providing a constant current of 35mA (5V at the monitor out).
After the debugging, the UL channel became the least noisy among the OSEM channels! Mode cleaner was able to lock and maintain it.
We came in the morning with the following scene on the zita monitor:
The MC1 watchdog was tripped and seemed like IMC struggled all night with misconfigured WFS offsets. After restoring the MC1 WD, clearing the WFS offsets, and seeing the suspension damp, the MC caught lock. It wasn't long before the MC unlocked, and the MC1 WD tripped again.
We tried few things, not sure what order we tried them in:
Nothing worked. We kept seeing that ULPD var on MC1 keeps showing kicks every few minutes which jolts the suspension loops. So we decided to record some data with PSL shutter closed and just suspension loops on. Then we switched off the loops and recorded some data with freely swinging optic. Even when optic was freely swinging, we could see impulses in the MC1 OSEM UL PD var which were completely uncorrelated with any seismic activity. Infact, last night was one fo teh calmer nights seismically speaking. See attachment 2 for the time series of OSEM PD variance. Red region is when the coil outputs were disabled.
Edit Thu May 13 14:47:25 2021 :
Added OSEM Sensor timeseries data on the plots as well. The UL OSEM sensor data is the only channel which is jumping hapazardly (even during free swinging time) and varying by +/- 30. Other sensors only show some noise around a stable position as should be the case for a freely suspended optic.
Here are the mass properties for the only the test mass assembly (optic, 3" ring, and wire block). (Updated with g*mm^2)
No, this is the property of the suspension assembly. The mass says 10kg
Could you do the same for the testmass assembly (only the suspended part)? The units are good, but I expect that the values will be small. I want to keep at least three significant digits.
Working with Chris, we decided that it is probably better to use a simple filter module as a controller before we make the model more complicated. I will use the plant model that I have already made (see attachment 1 of this). then attach a single control filter module to that: as seen in attachment 1. because I only want to work with one degree of freedom (position) I will average the four outputs which should give me the position. Then by feeding the same signal to all four inputs I should isolate one degree of freedom while still using the premade plant model.
The model I made that is shown in attachment 2 is the model I made from the plan. And it complies! yay! I think there is a better way to do the average than the way I showed. And since the model is feeding back on itself I think I need to add a delay which Rana noted a while ago. I think it was a UnitDelay (see page 41 of RTS Developer’s Guide). So I will add that if we run into problems but I think there is enough going on that it might already be delayed.
Since our model (x1sup_isolated.mdl) has compiled we can open the medm screens for it. I provide a procedure below which is based on Jon's post.
$ cd docker-cymac
$ eval $(./env_cymac)
$ medm -x /opt/rtcds/tst/x1/medm/x1sup_isolated/X1SUP_ISOLATED_GDS_TP.adl
To see a list of all medm screens use:
$ cd docker-cymac
# cd /opt/rtcds/tst/x1/medm/x1sup_isolated
Some of the other useful ones are:
See attachment 4. This screen shows the POS plant filter module that will be filled by the filter representing the transfer function of a damped harmonic oscillator:
THIS TF HAS BEEN UPDATED SEE NEXT POST
The first one of these screens that are of interest to us (shown in attachment 3) is the X1SUP_ISOLATED_GDS_TP.adl screen, which is the CDS runtime diagnostics screen. This screen tells us "the success/fail state of the model and all its dependencies." I am still figuring out these screens and the best guide is T1100625.
The next step is taking some data and seeing if I can see the position damp over time. To do this I need to:
We picked a few parameters from 40m summary page and plotted them to see the effect of new settings. On April 4th, old settings were present. On April 28th (16091), new input matrices and F2A filters were uploaded but suspension gains remained the same. On May 5th (16120), we uploaded new (higher) suspension gains. We chose Sundays on UTC so that it lies on weekends for us. Most probably nobody entered 40m and it was calmer in the institute as well.
We can download data and plot comparisons ourselves and maybe calculate the spectrums of MC_TRANS_PIT/YAW and MC_REFL_DC when IMC was locked. But we want to know if anyone has better ways of characterizing the settings that we should know of before we get into this large data handling which might be time-consuming. From this preliminary 40m summary page plots, maybe it is already clear that we should go back to old settings. Awaiting orders.
Attached is the control loop diagram when main laser is locked to IMC and a single arm (XARM) is locked to the transmitted light from IMC.
Did you match the local PC time with the GPS time?
The next step is to provide the 65 kHz clock signals from the timing fanout via LC optical fiber. I overlooked the fact that an SPX optical transceiver is required to interface the fiber to the timing slave board. These were not provided with the timing slaves we received. The timing slaves require a particular type of transceiver, 100base-FX/OC-3, which we did not have on hand. (For future reference, there is a handy list of compatible transceivers in E080541, p. 14.) I placed a Digikey order for two Finisar FTLF1217P2BTL, which should arrive within two days.
Today I brought and installed the new optical transceivers (Finisar FTLF1217P2BTL) for the two timing slaves. The timing slaves appear to phase-lock to the clocking signal from the master fanout. A few seconds after each timing slave is powered on, its status LED begins steadily blinking at 1 Hz, just as in the existing 40m systems.
However, some other timing issue remains unresolved. When the IOP model is started (on either FE), the DACKILL watchdog appears to start in a tripped state. Then after a few minutes of running, the TIM and ADC indicators go down as well. This makes me suspect the sample clocks are not really phase-locked. However, the models do start up with no error messages. Will continue to debug...
A few corrections to last analysis:
We today measured the calibration factors for XARM_OUT and YARM_OUT in nm/cts and replotted our results from 16117 with the correct frequency dependence.
Calibration of XARM_OUT and YARM_OUT
Inferring noise contributions to arm cavities:
Edit Mon May 10 18:31:52 2021
See corrections in 16129.
I copied c1scx.mdl to the docker to attach to the plant using the commands:
$ ssh nodus.ligo.caltech.edu
$ cd opt/rtcds/userapps/release/isc/c1/models/simPlant
$ scp c1scx.mdl controls@c1sim:/home/controls/docker-cymac/userapps
Here's my first attempt at doing angular actuation calibration for IMC mirrors using the method descibed in /users/OLD/kakeru/oplev_calibration/oplev.pdf by Kakeru Takahashi. The key is to see how much is the cavity mode misaligned from the input mode of beam as the mirrors are moved along PIT or YAW.
There two possible kinds of mismatch:
Kakeru's document goes through cases for linear cavities. For IMC, the mode mismatches are bit different. Here's my take on them:
Calibration factor at DC [µrad/cts]
When using mdl2adl I was getting the error:
$ cd /home/controls/mdl2adl
$ ./mdl2adl x1sup.mdl
error: set $site and $ifo environment variables
to set these in the terminal use the following commands:
$ export site=tst
$ export ifo=x1
On most of the systems, there is a script that automatically runs when a terminal is opened that sets these but that hasn't been added here so you must run these commands every time you open the terminal when you are using mdl2adl.
I added the IPC parts back to the plant model so that should be done now. It looks like this again here.
I can't seem to find the control model which should look like this. When I open sus_single_control.mdl, it just shows the C1_SUS_SINGLE_PLANT.mdl model. Which should not be the case.
Assembled chassis from De Leone placed in the 40 Meter Lab, along the west wall and under the display pedestal table. The leftover parts are in smaller Really Useful boxes, also on the parts pile along the west wall.
We have uploaded the new damping gains on all the suspensions of IMC. This completes changing all the configuration to as mentioned in 16066 and 16072. The old setting can be restored by running python3 /users/anchal/20210505_IMC_Tuned_SUS_with_Gains/restoreOldConfigIMC.py from allegra or donatella.
I put the box containing the untested OSEMs from KAGRA near the south flow bench on the floor.